# A Survey of Deep Learning-based Radiology Report Generation Using Multimodal Data

### v1

@@ 3.5.2 Retrieval-based and template-based techniques
3.5.3 Large language model to assist report generation
With the emergence of ChatGPT (OpenAI, 2023), its powerful language abilities made researchers eager to harness the power of large language models to aid in report generation. However, employing it directly within the medical domain led to unsatisfactory outcomes (Tu et al., 2024; Yan et al., 2023; Sun et al., 2023). Two works (Selivanov et al., 2023; Wang et al., 2023a) initially predicted report-related information such as diseases, lesion regions, and visual features, and generated preliminary reports. Subsequently, they employed pre-trained large language models, e.g., ChatGPT (Wang et al., 2023a) or GPT-3 (Brown et al., 2020; Selivanov et al., 2023; Wang et al., 2023a) to improve the preliminary report with the predicted information and produced the final reports. The results have been moderately improved. Exploration of large language models in the field of medical report generation still requires further investigation, which is discussed in Section 7.4.
3.6 Training strategy
Training strategy refers to the techniques used to train neural network models. Traditionally, models are trained by minimizing various loss functions. Therefore, this section begins by introducing different loss functions (Section 3.6.1), followed by a discussion on reinforcement learning (Section 3.6.2), and the curriculum learning’s application to the report generation task (Section 3.6.3).
3.6.1 Loss functions
The application of an auxiliary loss function can provide additional supervision signals, further enhancing model performance. Two works Wang et al. (2021); Li et al. (2022a) applied an additional constraint between features extracted from the generated and real reports. Zhang et al. (2023b) created two different versions of an input image by adding noise and feeding them into two networks. An auxiliary loss function ensures the consistency of the outputs produced by the two generators. Wang et al. (2024a) obtained two discriminative regions in an image from the generated words and visual classifier separately, then enforced the consistency between them.
3.6.2 Reinforcement learning
Reinforcement learning involves training an agent to make optimal decisions through trial and error, aiming to maximize targeting rewards. It offers a method to update the model parameters based on non-differentiable reward functions (Messina et al., 2022). Evaluation metrics can be considered as rewards, such as CIDEr (Kaur and Mittal, 2022b), BLEU (Qin and Song, 2022; Gu et al., 2024), METEOR (Qin and Song, 2022), ROUGE (Qin and Song, 2022), BERTScore (Miura et al., 2021), F1 score (Miura et al., 2021), and accuracy (Hou et al., 2021b). In addition, Hou et al. (2021b) trained a language fluency discriminator using the ground truth and generated reports, and then utilized the discriminator to provide rewards.
3.6.3 Curriculum learning
Liu et al. (2021a) utilized curriculum learning (Platanios et al., 2019) to classify training instances and trained a model from simple to complex samples. Data pairs were evaluated based on image heuristics, image confidence, text heuristics, and text confidence. Image heuristic evaluated the similarity between input images and normal images. Image confidence indicated the confidence of a classification model. The report heuristic was related to the number of abnormal sentences, and report confidence was evaluated by the negative log-likelihood loss (Xu et al., 2020).
@@ 4 Datasets
Datasets play a crucial role in the development of report-generation models. Abundant and diverse training data can improve the model’s accuracy and generalizability. Moreover, a suitable test dataset makes it realistic to test the model’s performance in a practical scenario. In this section, we selected 11 public medical image-report datasets utilized in the reviewed articles to provide a comprehensive introduction of popular and newly collected datasets, see Table 1. The usage of datasets in each article is shown in Table LABEL:tab:overall in Appendix A. The datasets primarily focus on the lung, including X-ray (Demner-Fushman et al., 2016; Johnson et al., 2019a, b) and CT (Li et al., 2020; Liu et al., 2021e). In addition, there are also publicly available datasets on eye scans (Huang et al., 2021b; Lin et al., 2023; Li et al., 2021) and breast scans (Yang et al., 2021a). The concentration of medical report generation efforts on chest X-rays can be attributed, in part, to the accessibility of large-scale publicly available datasets.
@@
The most popular datasets are the Indiana University Chest X-Ray Collection (IU X-Ray) (Demner-Fushman et al., 2016) and the MIMIC Chest X-ray (MIMIC-CXR) (Johnson et al., 2019a, b). The IU X-ray contains 7470 images of frontal and lateral X-rays and 3955 reports with manual annotation based on the MeSH codes [^3] https://www.nlm.nih.gov/mesh/meshhome.html and the RadLex codes. These codes encompass standard medical terminology, such as anatomical structures, diseases, pathological signs, foreign objects, and attributes, as defined by authoritative institutions. For deep learning methods, the size of the IU X-RAY is insufficient, while the collection of MIMIC-CXR alleviates this problem (Messina et al., 2022). It consists of 377,110 images and 227,827 reports, and has been multi-labeled by automatic tools according to 14 disease categories. An image-report sample from the MIMIC-CXR dataset is shown in the Figure 4. The indication, technique, and comparison provide fundamental information for a Chest X-ray test. The automation of radiology report generation targets the findings and impressions sections. The findings section provides a detailed description of the entire image, while the impressions section summarizes these observations. The follow-up work provides MIMIC-CXR with richer label information. The Chest ImaGenome dataset (Wu et al., 2021) is based on the anteroposterior and posteroanterior view Chest X-ray images in the MIMIC-CXR dataset. It provides an anatomy-centered scene graph for each image, including anatomical location and relation annotations. The annotations are generated through two automated pipelines, and a separate dataset of 500 manually annotated cases is also provided for testing.
@@
Increasing attention has been paid to the collection of ophthalmic image-report pairs (Huang et al., 2021b; Li et al., 2021; Lin et al., 2023). Ophthalmic image has lots of modalities, such as Fluorescein Angiography (FA), Fundus Fluorescein Angiography (FFA), Color Fundus photography (CFP), fundus photograph (FP), optical coherence tomography (OCT), fundus autofluorescence (FAF), Indocyanine Green Chorioangiography (ICG), and red-free filtered fundus images. Most of retinal datasets are based on one or two modalities and the last released Retina ImBank dataset is the first multi-modality retinal image-text dataset (Lin et al., 2023). Some datasets provide additional labels for each image, such as lesion boundary (Li et al., 2021), lesion category (Huang et al., 2021b; Li et al., 2021), and keywords (Huang et al., 2021b).

### v2

@@ 3.5.2 Retrieval-based and template-based techniques for report generation
In conclusion, retrieval-based and template-based techniques utilize existing data or predefined templates, theoretically enhancing performance, particularly in linguistic fluency. However, challenges such as inadequate feature representation and immature similarity measurement often lead to inferior outcomes. For instance, while inadequate feature representation may only affect specific generated words in decoder-based models, it can result in entirely irrelevant sentence selections in retrieval-based and template-based approaches. Furthermore, these techniques generally exhibit lower flexibility and limited content generation, negatively impacting human-computer interaction.
4 Large Models in Report Generation
Large models, including language models, vision models and multi-modal models, have revolutionized the field of deep learning. Particularly for large language models, their human-like conversational capabilities and ability to respond to free-text queries without task-specific training have been profoundly impactful. However, deploying general large language models in clinical settings remains unfeasible due to insufficient response accuracy and safety concerns (Yan et al., 2023; Saab et al., 2024), which has led to significant attention on their application to medical tasks such as report generation.
In the initial phases, rather than directly generating reports from imaging data, researchers focus on using general large language models to revise reports produced by traditional methods (Selivanov et al., 2023; Wang et al., 2023a). Traditional deep learning methods were applied to predict report-related information (e.g., lesion regions) and generate preliminary reports from images. Subsequently, pre-trained models (e.g., ChatGPT (Wang et al., 2023a) and GPT-3 (Brown et al., 2020; Selivanov et al., 2023; Wang et al., 2023a)) were employed to improve these reports based on the predicted information, leading to moderate improvements.
Training a large model on a large medical dataset is a more straightforward solution to achieve better performance. An overview of these models is presented in Table 1. We have provided the GitHub repository link for the open-source model. Some links provide only inference APIs (Alkhaldi et al., 2024; Chen et al., 2024), while others include training code (Zhang et al., 2024a; Li et al., 2024; Wang et al., 2023e; Pellegrini et al., 2023b; Wu et al., 2023a). This process generally comprises two main phases. The first phase is pre-training, where the model learns general feature representations through unsupervised or self-supervised learning on a large, unlabeled medical dataset. This phase enables the model to acquire broad generalization capabilities and adjust its weights to be more aligned with the target medical tasks. The second phase is fine-tuning, which involves refining the model for a specific task (e.g., medical report generation) using a relatively small, high-quality labeled dataset. The common monitoring tool is wandb [^3] https://wandb.ai/home, tracking metrics like loss or validation set results, similar to traditional training. Understanding this process requires grasping three key components:
Table 1: Comparison of vision–language models for medical tasks. “Params" describes the parameter scale, roughly indicating the size of LLMs. “Foud" indicates if the model can be used for other medical tasks besides report generation. “RG" shows if it is trained or tested for report generation. “OS" refers to its open-source status.
Name Params Visual Encoder Adapter LLM Foud RG OS
RaDialog
(Pellegrini et al., 2023b)
7B/13B/33B BioViL-T Q-Former Vicuna-7B/13B/33B https://github.com/ChantalMP/RaDialog
R2GenGPT
(Wang et al., 2023e)
7B Swin Transformer Linear LLaMA2-7B https://github.com/wang-zhanyu/R2GenGPT
LLM-CXR
(Lee et al., 2023)
13B VQ-GAN LLaMA2-13B
MAIRA-1
(Hyland et al., 2023)
7B RAD-DINO MLP-4 Vicuna-7B
MAIRA-2
(Bannur et al., 2024)
7B/13B RAD-DINO MLP-4 Vicuna-7B/13B v1.5
CheXagent
(Chen et al., 2024)
7B EVA-CLIP Q-Former Mistral-7B-v0.1 https://github.com/Stanford-AIMI/CheXagent
LLava-Med
(Li et al., 2024)
7B BioMedCLIP Linear LLaMA-7B https://github.com/microsoft/LLaVA-Med
MiniGPT-Med
(Alkhaldi et al., 2024)
7B EVA-CLIP Linear LLaMA2-7B https://github.com/Vision-CAIR/MiniGPT-Med
LLaVA-Ultra
(Guo et al., 2024)
13B CLIP-ViT-L/14 and SAM-ViT-L Linear LLaMA-13B
BiomedGPT-B
(Zhang et al., 2024a)
182M ResNet-101 BERT-style encoder
+ GPT-style decoder
https://github.com/taokz/BiomedGPT
RadFM
(Wu et al., 2023a)
14B 3D ViT Transformer +linear MedLLaMA-13B https://github.com/chaoyi-wu/RadFM
MedVersa
(Zhou et al., 2024)
Swin Transformer /3D UNet Linear Llama-2
  • Acquisition of medical image-text datasets with prompts designed for training.
  • Except for BiomedGPT (Zhang et al., 2024a), the other reviewed papers are based on large language models. Thus, understanding the transition from large language models to large multimodal models, particularly regarding the integration of image data, is essential.
  • Training large models demands substantial computational resources. For instance, Touvron et al. (2023) reported that training the famous large language model named LLaMA with 65B-parameter required 2,048 A100 GPUs with 80GB of RAM over approximately 21 days. Such resource demands are beyond the reach of most researchers. Therefore, it is crucial to explore methods for training large medical models with limited resources.
We discussed the aforementioned issues in Sections 4.1, 4.2, and 4.3, respectively. It is untenable to fully separate large models for generating medical reports from foundational medical models. Thus, in these sections, foundational model methods are introduced first, followed by their application in report generation.
4.1 Acquisition of medical image-text datasets with prompts
A large dataset is required during the pre-training phase. PMC-15M (Zhang et al., 2023c) is a dataset comprising 15 million biomedical image-text pairs, each consisting of an image and its corresponding caption from biomedical papers. In addition to medical images, it also includes other types of data, such as general biomedical illustrations and microscopy images, without any prompts provided. Therefore, it is impractical to directly utilize the entire dataset for medical image analysis. Li et al. (2024) derived two sub-datasets from PMC-15M for large model training. The first dataset, consisting of 600K image-text pairs, was selected based on caption length. Corresponding prompts were randomly chosen from a predefined list, such as “Describe the image concisely”. The second dataset consists of multi-round questions and answers for 60K filtered images from the PMC-15M dataset. The image types include X-rays, CT scans, MRIs, histopathology slides, and gross pathology images. In addition to the original text associated with each image, sentences referring to these images were extracted from the original papers to provide extra context. Image, original text, and additional descriptions were then provided to GPT-4, which was tasked with generating multi-round questions and answers as if it could see the images. This novel dataset design enables the Large Language and Vision Assistant for BioMedicine (LLaVa-Med) (Li et al., 2024) to perform well through pre-training on it and also inspired subsequent works (Guo et al., 2024).
A 60k medical image dataset is relatively small for large model training, and the image-text pair format limits its utility for diverse medical tasks such as disease diagnosis. Consequently, efforts have been made to compile a comprehensive radiology-focused dataset, with the most straightforward approach being the integration of multiple public datasets (Alkhaldi et al., 2024; Zhang et al., 2024a; Chen et al., 2024; Wu et al., 2023a). For example, Zhang et al. (2024a) aggregated 14 public datasets and Chen et al. (2024) derived a large-scale instruction-tuning dataset about Chest X-rays from 28 publicly available datasets. Distinct prompt templates were designed for each type of task. For report generation, most studies incorporate MIMIC-CXR into the training set (Alkhaldi et al., 2024), with a few also utilizing the Indiana University Chest X-Ray Collection (IU-Xray) (Zhang et al., 2024a). In addition, most datasets contained only 2D images. Wu et al. (2023a) integrated 18 public datasets to construct a radiology image dataset called the Medical Multi-modal Dataset (MedMD), which includes 15.5M 2D scans and 500k 3D scans. They also filtered MedMD, retaining only radiology images. However, no studies have yet conducted ablation experiments to assess how pre-training on different datasets affects model performance. For example, whether using the MIMIC-CXR dataset in both the pre-training and fine-tuning phases improves report generation compared to using it only in fine-tuning.
For the fine-tuning phase, prompts need to be designed based on the existing report generation dataset (e.g., MIMIC-CXR). Most reviewed papers rely on fixed prompts, such as “Generate a comprehensive and detailed diagnosis report for this chest X-ray image” (Wang et al., 2023e), or use prompts based on the other sections in the report (i.e., indication, technique, and comparison) to provide the model with additional patient-related information (Hyland et al., 2023; Bannur et al., 2024). Pellegrini et al. (2023b) integrated disease classification results into the prompt. However, the classifier’s low accuracy yields minimal improvement. Given that prompt design is critical for optimizing model performance (Liu et al., 2023b), further research into developing effective prompts for report-generation tasks is necessary.
4.2 Evolution from large language models to large multimodal models
A common approach for evolving from a large language model to a large multimodal model is to add an adapter between the visual encoder and the language model to align image features with the language model’s input, as shown in Figure 6. Table 1 lists the adapter designs, which are usually linear projections or multilayer perceptrons (MLP). Hyland et al. (2023) found that a 4-layer MLP outperforms a 2-layer MLP on report generation tasks, due to the increased non-linearity enhancing the model’s expressive power. Additionally, Pellegrini et al. (2023b) introduced a more sophisticated adapter (i.e., Q-Former) from Bilp-2 (Li et al., 2023a), but it yielded no significant improvement.
When training large foundational medical models, task-specific prompts are designed, such as “Does the patient have disease X?" for disease diagnosis, “Is this image from modality X?" for modality recognition, and “Generate a detailed diagnosis report for this chest x-ray image." for report generation. Typically, cross-entropy loss is applied between the predicted and ground-truth text sequences, helping to standardize the training process across different tasks.
Figure 6: A typical process for transitioning from language models to multimodal models.
Based on the architecture, the evolution can be achieved in one step by training the visual encoder, adapter, or large language model directly with the target dataset (Wang et al., 2023e; Hyland et al., 2023). The model’s initial weights come from general large language models (e.g., Vicuna (Chiang et al., 2023), LLaMA (Touvron et al., 2023), Mistral (Jiang et al., 2023)), general visual encoder (e.g., EVA-CLIP (Sun et al., 2023)), and domain-specific visual encoders (e.g., BioMedCLIP (Zhang et al., 2023c), BioViL(Bannur et al., 2023), RAD-DINO (Pérez-García et al., 2024)). Wang et al. (2023e) found that for report generation, training only the adapter was ineffective compared to training both the adapter and the visual encoder. Hyland et al. (2023) observed that when the visual encoder is frozen and only the adapter and LLM are trained, the choice of a radiology-specific visual encoder can significantly improve model performance. Guo et al. (2024) argued that the size imbalance between the visual encoder and the language model causes a bottleneck in feature extraction. To address this, they use two large visual encoders, CLIP and the Segment Anything Model (SAM) (Kirillov et al., 2023), and combine their features through weighted addition, which outperforms the use of a single encoder.
The one-step approach is highly convenient; however, training the entire visual encoder or LLM on large datasets is costly. To address this, two-step approaches have been proposed for scenarios that require training on larger datasets to achieve better generalization. Inspired by LLaVA (Liu et al., 2024), LLaVa-Med (Li et al., 2024) presents a feasible solution. In the initial phase, the 600K dataset from PMC-15M is used to train the adapter for image captioning, with the visual encoder and language model parameters kept frozen. In the subsequent phase, the model is trained on the 60K instruction set from PMC-15M. The visual encoder weights remain fixed, and the projection layer and language model parameters are updated.
Lee et al. (2023) argued that using adapters hinders the interaction between visual and semantic information. To address this, they proposed using frozen Vector Quantised Variational Autoencoder (VQ-GAN) (Esser et al., 2021) to represent image information in a discrete form, incorporating these image tokens as special tokens into the vocabulary of large language models. While experimental results have yet to show improvements in report generation tasks using this method compared to adapters, the VQ-GAN approach remains notable. It introduces a learnable codebook that maps encoded image features to the nearest discrete variable, enabling discretization. This technique is well-established in natural image generation for capturing high-quality discrete image representations but remains underexplored in report generation.
Most medical foundation models are trained directly on multi-task datasets, enabling them to perform a variety of tasks. In contrast, Zhou et al. (2024) treat the large model as a planner, predicting specific tokens that activate the corresponding vision modules (e.g., detection and segmentation), which allows it to handle an even wider array of tasks. For instance, when the <DET> token is predicted, the detection module is triggered to generate bounding box information. However, this approach increases the number of models involved, making the pipeline non-end-to-end.
Integrating image features into large language models while preserving their conversational abilities is crucial. However, the current report-generation large models do not achieve this. General medical models are only capable of handling simple conversations. Nonetheless, as shown in Section 8, with accessing to sufficient data, large models demonstrate advantages on natural language-based evaluation metric, highlighting their potential.
4.3 Training large medical models with limited resources
Directly applying a general-purpose LLM into a highly specialized task (e.g. radiology report generation) can not achieve an optimal performance, therefore normally requires LLM fine-tuning using a task-specific dataset. However, fine-tuning the entire LLM is extremely resources demanding. Employing adapters, frozen layers, and training only a subset of the LLM parameters can conserve training resources; however, these techniques often involve a trade-off between efficiency and model performance. Additionally, incorporating extra adapters increases the overall model parameters, further burdening the computational resources required during the inference stage and prolonging the inference time. This section presents two more direct approaches for conserving training resources: Low-Rank Adaptation (LoRA) (Hu et al., 2021), a classic large model fine-tuning technique widely utilized in report generation task (Wang et al., 2023e; Alkhaldi et al., 2024; Pellegrini et al., 2023b; Guo et al., 2024), and a lightweight medical foundation model for medical tasks, including report generation.
When fine-tuning large models, the weights $W$ update can be expressed as $W+\Delta W$, where $\Delta W$ denotes the change in $W$ during training. The core idea of LoRA  [^4] https://github.com/microsoft/LoRA is that not all parameters need to be changed for a specific target task (e.g., report generation) during fine-tuning. The update weights $\Delta W$ can be efficiently represented using a low-rank matrix and computed via low-rank decomposition. Specifically, given model weights $W\in\mathbb{R}^{d\times k}$, the weight update is assumed to be of low rank $r$, i.e., $W+\Delta W=W+BA$, where $B\in\mathbb{R}^{d\times r}$ and $A\in\mathbb{R}^{r\times k}$, with $r\ll\min(d,k)$. Thus, only the low-rank matrices $A$ and $B$ are trained, and their product $BA$ is added to the original weights. This approach substantially reduces the computational resources required, as only the delta weights $\Delta W$ need to be stored for each new model, thereby conserving disk space. Moreover, experiments conducted on large models (such as GPT-3 175B) have demonstrated that models fine-tuned with LoRA can achieve performance comparable to, or even surpassing, that of full model fine-tuning. However, Wang et al. (2023e) observed that when applying this approach to fine-tune the adapter and visual encoder (with 90.9M parameters) for report generation, the performance was inferior to that of full model fine-tuning. This discrepancy arises because, as the model size decreases, the proportion of effective parameters increases, thereby diminishing the advantage of using low-rank matrices to approximate $\Delta W$. The LoRA has been integrated into the State-of-the-art Parameter-Efficient Fine-Tuning (PEFT) library  [^5] https://github.com/huggingface/peft and can be directly accessed via its interface. PEFT is designed to efficiently fine-tune large pre-trained models.
A straightforward strategy to conserve training resources is to reduce the model’s trainable parameters. However, due to the performance improvements (Bannur et al., 2024) and enhanced generalization ability with larger models, most reviewed studies still employ models exceeding 7B parameters (see Table 1). Recently, Zhang et al. (2024a) developed lightweight medical foundation models of 33M/93M/182M parameters, with the 182M BiomedGPT-B achieving state-of-the-art results in 16 out of 25 medical tasks compared to 7B models. Although it has not been compared with other models on the standard report generation benchmark, it offers a valuable insight that relatively lightweight large models can be trained to accomplish report generation. BiomedGPT mainly differs from traditional models by curating a large-scale pretraining corpus and specifically developing instruction-tuning data, emphasizing the importance of a high-quality labeled large training dataset. Although for the report generation task, BiomedGPT-B was tested on only 30 image-report pairs and evaluated manually rather than using natural language or medical correctness metrics, it is still regarded as a significant advancement for lightweight foundation models in medical report generation.
It is noteworthy that there remains no definitive consensus on the optimal model size and data quantity requirements for training large models for report generation tasks. Pellegrini et al. (2023b) fine-tuned models with 7B, 13B, and 33B parameters on the MIMIC dataset, finding no significant performance differences. In contrast, Bannur et al. (2024) merged four datasets as training datasets, introducing more diverse inputs and outputs rich in information (as described in Section 5.2). They trained both 7B and 13B models, observing performance improvements as model size increased (see Table 4). Similarly, Zhang et al. (2024a) noted consistent performance enhancements with models of 33M, 92M, and 182M parameters. Despite differences in architecture, these findings suggest a bottleneck effect of model size and data usage on model performance, underscoring the need for systematic investigation.
5 Explainability
Explainability of a deep learning model is defined as the ability to justify an outcome in understandable terms for a human (Doshi-Velez and Kim, 2017; Messina et al., 2022). This is especially critical in medical contexts, where model outputs can directly or indirectly influence clinical decision-making. However, most current research focuses on improving report generation model performance, with relatively few studies considering explainability. This section summarizes recent advancements in this area, including visualization, auxiliary tasks, and uncertainty quantification.
5.1 Visualization for explainability
Aligned with previous surveys (Messina et al., 2022), visualization remains the dominant method for interpreting model outputs. It highlights the most crucial part of the inputs relevant to report generation, such as specific areas in images, keywords in text, and certain nodes in knowledge graphs. Visualization allows radiologists to assess whether the model accurately targets the relevant information, thereby fostering greater confidence in the system. The reviewed report generation methods mainly employ two approaches to performing visualization.
The first approach is a gradient-based method named Gradient-weighted Class Activation Mapping (Grad-CAM) (Selvaraju et al., 2017). It calculates the importance of each image feature map for a particular prediction by weighting them with the gradients and then performing a weighted sum. Report generation methods often use Grad-CAM in auxiliary classification tasks to identify image regions that influence model decisions (Alfarghaly et al., 2021; Huang et al., 2021b; Pino et al., 2021; Kaur and Mittal, 2022b), with some methods applying it directly to report generation tasks to determine the image regions relevant to report generation (Najdenkoska et al., 2021, 2022).
The second approach utilizes the attention weights. Compared to Grad-CAM, which is limited to images, attention-based visualization is more applicable across various input data modalities, including images (Liu et al., 2021b; Chen et al., 2021; Wang et al., 2023d; Zhang et al., 2023b; Li et al., 2022a; Qin and Song, 2022; Zhang et al., 2023a; Moon et al., 2022; Hou et al., 2021a; Cao et al., 2023; Lin et al., 2023; You et al., 2022; Li et al., 2023e; Kong et al., 2022; Xue et al., 2024), medical terminology (Liu et al., 2021b), retrieved real reports (Liu et al., 2021b), and medical knowledge base (Liu et al., 2021b). Typically, attention weights are derived from the Scaled Dot-Product Attention mechanism (Vaswani et al., 2017), which accepts query (Q), key (K), and value (V) vectors as inputs. The formulation is expressed as $Attention(Q,K,V)=softmax(\frac{QK^{T}}{\sqrt{d_{k}}})V$, where $\sqrt{d_{k}}$ is a scaling factor related to the input dimension. The expression $softmax(\frac{QK^{T}}{\sqrt{d_{k}}})$ is interpreted as the attention weights.
Regardless of visualization methods, studies typically provide limited image examples to illustrate that their models can accurately identify relevant regions; however, they are deficient in the systematic evaluation of large-scale datasets. Moreover, there is a lack of quantitative metrics to assess visualization results, possibly due to the absence of ground truth. Non-image-based ground truth can be obtained through text comparisons, while image-based ground truth can be established by collecting physicians’ eye-tracking data or using bounding boxes for lesions or anatomical structures.
5.2 Auxiliary tasks for explainability
Auxiliary tasks can provide valuable information for image classification or detection (see Section 3.3.1), thus enhancing system interpretability by offering radiologists additional insights. However, the potential of auxiliary tasks in improving interpretability has been largely underutilized in previous works (Messina et al., 2022). Tanida et al. (2023) addressed this gap by linking output results to specific image regions via object detection, enabling users to select areas of interest and receive corresponding explanations. Dalla Serra et al. (2023b) introduced a multi-label classification head that classifies findings for each detected anatomical region. This model outputs both the bounding box locations and findings as intermediate results, which increases model transparency. Bannur et al. (2024) outputs the corresponding bounding box while generating each lesion-containing description.
5.3 Uncertainty quantification for explainability
The uncertainty discussed in this section quantifies the model’s confidence in the quality of the generated outputs. Wang et al. (2024b) measured the uncertainty based on Monte Carlo dropout variational inference (Gal and Ghahramani, 2016). When dropout is applied in the network, performing $T$ stochastic forward passes for a given input allows the variance of the outputs to reflect the model’s uncertainty. Greater variance implies greater uncertainty. They defined report-level textual uncertainty and sentence-level textual uncertainty. Report-level uncertainty is measured by the variance of similarities between the $T$ generated reports. For sentence-level uncertainty, they first identify the report with the minimum similarity distance to other reports from the $T$ outputs, termed the average reference report. The uncertainty of each sentence is then represented by the variance of the similarities between the sentences in the average reference report and their corresponding sentences in the $T$ generated reports. Compared to report-level uncertainty, sentence-level uncertainty enables physicians and patients to assign varying trust levels to individual sentences rather than fully affirming or rejecting the entire report, warranting further study.
@@ 6 Datasets
Datasets play a crucial role in the development of report-generation models. Abundant and diverse training data can improve the model’s accuracy and generalizability. Moreover, a suitable test dataset makes it realistic to test the model’s performance in a practical scenario. In this section, we selected 11 public medical image-report datasets utilized in the reviewed articles to provide a comprehensive introduction to popular and newly collected datasets, see Table 2. The usage of datasets in each article is shown in Table LABEL:tab:overall in Appendix A. The datasets primarily focus on the lung, including X-ray (Demner-Fushman et al., 2016; Johnson et al., 2019a, b) and CT (Li et al., 2020; Liu et al., 2021e). In addition, there are also publicly available datasets on eye scans (Huang et al., 2021b; Lin et al., 2023; Li et al., 2021) and breast scans (Yang et al., 2021a). The concentration of medical report generation efforts on chest X-rays can be attributed, in part, to the accessibility of large-scale publicly available datasets.
@@
The most popular datasets are the IU-Xray (Demner-Fushman et al., 2016) and the MIMIC-CXR (Johnson et al., 2019a, b). The IU-Xray contains 7470 images of frontal and lateral X-rays and 3955 reports with manual annotation based on the MeSH codes [^6] https://www.nlm.nih.gov/mesh/meshhome.html and the RadLex codes. These codes encompass standard medical terminology, such as anatomical structures, diseases, pathological signs, foreign objects, and attributes, as defined by authoritative institutions. Manual keyword annotation is invaluable because it helps standardize synonym expressions and clarify linguistic ambiguities, which facilitates the structuring of information in free-text reports (Pellegrini et al., 2023a). However, IU-Xray annotations are designed to improve retrieval efficiency rather than to structure information, which may result in the overlooking of speculative details when converting free-text reports into a structured format. For example, the annotation for “Focal airspace disease in the right middle lobe. This is most concerning for pneumonia" is annotated as “Airspace Disease/lung/middle lobe/right/focal", which omits the concern for pneumonia. Additionally, the IU-Xray dataset does not provide an official split strategy. Most reviewed studies use a 7:1:2 ratio for training, validation, and test sets. For deep learning methods, the size of the IU-Xray is insufficient, while the collection of MIMIC-CXR alleviates this problem (Messina et al., 2022). Chen et al. (2020) also noted that compared to IU-Xray, MIMIC-CXR may exhibit a greater diversity in report patterns. This is more conducive to training a model with strong generalization performance. MIMIC-CXR consists of 377,110 images and 227,827 reports and has been multi-labeled by automatic tools into 14 categories, including lung-related diseases (e.g., pneumonia), heart-related diseases (e.g., cardiomegaly), pleural-related diseases (e.g., pleural effusion), bone-related conditions (e.g., fracture), support devices, and cases where no abnormal finding is detected. The F1 scores for consolidation, pneumonia, atelectasis, pneumothorax, pleural effusion, and fracture annotation all exceed 0.9, indicating the reliability of these automatically annotated labels. An image-report sample from the MIMIC-CXR dataset is shown in Figure 7. The indication, technique, and comparison provide fundamental information for a Chest X-ray test. The automation of radiology report generation targets the findings and impressions sections. The findings section provides a detailed description of the entire image, while the impressions section summarizes these observations. The MIMIC-CXR dataset provides an official data split strategy, consisting of 368,960 training images with 222,758 reports, 2,991 validation images with 1,808 reports, and 5,159 test images with 3,269 reports. To facilitate different methods, many studies have preprocessed this dataset. The most well-known version, introduced by R2Gen (Chen et al., 2020), consists of 270,790 training images, 2,130 validation images, and 3,858 test images, and has been widely used in subsequent approaches. However, Chen et al. (2020) paired each report with only one image, treating multiple images for the same report (e.g., frontal and lateral views) as distinct cases. This can introduce errors, particularly in contrastive learning, as different views might be seen as negative pairs. Moreover, if a report references findings only visible in the lateral view but is matched with a frontal image, this misalignment could negatively impact the model training. Thus, while this data version is popular, careful selection is necessary when using it. Research on the role of frontal and lateral views in automated report generation has been a focus of previous studies. For instance, Bertrand et al. (2019) found that the lateral view alone is less effective than the frontal view, which provides more comprehensive information. However, combining both views can enhance the diagnosis of certain diseases, as they offer complementary data (Laserson et al., 2018). As discussed in Section 3.3.1.1, Wang et al. (2023c) demonstrated how this combination can be achieved using transformer attention. Nonetheless, utilizing both views increases computational costs and demands more processing power from devices.
The follow-up work provides MIMIC-CXR with richer label information. The Chest ImaGenome dataset (Wu et al., 2021a) is based on the anteroposterior and posteroanterior view Chest X-ray images in the MIMIC-CXR dataset. It provides an anatomy-centered scene graph for each image based on sentence-level annotation. The graph includes objects (e.g., lung), attributes (e.g., opacity), and their relationships. Bounding box locations are specified for key anatomical structures. The annotations are generated through two automated pipelines, and a separate dataset of 500 manually annotated cases is also provided for testing. Additionally, using timestamps from MIMIC-CXR, anatomical annotations are compared across sequential exams based on descriptive changes (e.g., decrease). Dalla Serra et al. (2023a) incorporated the anatomy information from prior reports into model training, improving most metrics. It should be noted that the ImaGenome dataset uses only frontal images, likely because lateral images have less influence on automatic report generation. Frontal images are generally sufficient for annotating most anatomical structures, and even in the presence of partial obstructions (e.g., up to 15% of the lung obscured by the heart or diaphragm (Raoof et al., 2012)), the annotation boxes remain largely unaffected. A new dataset, CheXpert Plus, comparable to MIMIC-CXR, was recently released by Stanford University (Chambon et al., 2024), containing 223,462 images and 187,711 reports. The studies for each patient are also arranged in chronological order. As an extension of the widely used chest X-ray classification dataset named CheXpert (Irvin et al., 2019), it adds corresponding reports, demographic information, automated labels for 14 disease categories, automated graph-based annotations, and a collection of pre-trained models built on this dataset. Although CheXpert Plus has fewer reports than MIMIC, the reports are longer, particularly in the impression section. CheXpert Plus is the largest available dataset for tasks related to the impression section. Additionally, it has been de-identified without altering the structure of the reports, making them more representative of real-world radiology text.
The above datasets contain X-ray images. Besides X-rays, CT is another important imaging modality for the chest. Publicly available datasets for CT report generation were primarily collected for COVID-19 research. The COV-CTR dataset (Li et al., 2020) includes 349 COVID and 379 Non-COVID images, sourced from published papers. One limitation of this dataset is its relatively low image quality. In contrast, the COVID-19 CT dataset (Liu et al., 2021e) contains higher-quality images, which were directly collected from hospitals, offering clearer diagnostic details.
@@
Increasing attention has been paid to the collection of ophthalmic image-report pairs (Huang et al., 2021b; Li et al., 2021; Lin et al., 2023). Ophthalmic image has lots of modalities, such as Fluorescein Angiography (FA), Fundus Fluorescein Angiography (FFA), Color Fundus photography (CFP), fundus photograph (FP), optical coherence tomography (OCT), fundus autofluorescence (FAF), Indocyanine Green Chorioangiography (ICG), and red-free filtered fundus images. Most retinal datasets are based on one or two modalities and the last released Retina ImBank dataset is the first multi-modality retinal image-text dataset (Lin et al., 2023). Some datasets provide additional labels for each image, such as lesion boundary (Li et al., 2021), lesion category (Huang et al., 2021b; Li et al., 2021), and keywords (Huang et al., 2021b).